OpenStudy (anonymous):

evaluate the improper integral or state that it diverges. integral of dx/(e^x + e^-x) from negative infinity to infinity.

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} \frac{ dx }{ e^{x}+e ^{-x} }\]

myininaya (myininaya):

I think it would help to get rid of that negative exponent by multiplying both top and bottom by e^x and then use a trig sub

OpenStudy (anonymous):

@myininaya i dont understand what you mean by using a trig substitution. doesnt that involve sin and cos not e??

OpenStudy (dumbcow):

\[\int\limits \frac{e^x}{e^{2x} +1}dx\] since \[\tan^2 +1 = \sec^2\] let \[e^x = \tan u\] \[e^x dx = \sec^2 u du\] \[\rightarrow \int\limits \frac{\sec^2 u}{\sec^2 u} du = \int\limits du = u + C\]

OpenStudy (dumbcow):

oops you dont need "+C" since you're evaluating at infinity

OpenStudy (anonymous):

but i need to find whether it diverges or converges using limits..

OpenStudy (dumbcow):

right \[\lim_{x \rightarrow \infty} \tan^{-1} (e^x) - \lim_{x \rightarrow -\infty} \tan^{-1} (e^x)\]